Week of 061211

Open Actions from last week:
  • follow up errors with GSSDATLAS (James/Simone) - In Progress

Chair: James

Gmod:

Smod:

Monday:

Log: LFC Problems over the weekend

New Actions:

  • Look at why sensors didn't pick up problems for LFC (James)

Discussion:

Tuesday:

Log: lfc-atlas local still a problem

New Actions:

  • check procedure for sys admin for log-db-error (James)
  • check why new sensor still doesn't work (James)

Discussion:

  • new sensor for LFC was put in place with timeout, but didn't seem to work - hard to test ;(

Wednesday

Log:

New Actions:

Discussion:

Thursday

Log:

New Actions:

Discussion:

Friday

Log:

New Actions:

Discussion:

Edit | Attach | Watch | Print version | History: r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r3 - 2007-02-02 - FlaviaDonno
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback